OpenStudy (anonymous):

Find the derivative of -2-6x. I have gotten to -2-6x-(-2-6x(x+h))/h

OpenStudy (solomonzelman):

the derivative of " f(x)-f(g) " would be f(x)' - f(g) ' the derivative of a -2 is zero (b/c it's a constant) and the derivative of 6x is \[x^n~'=nx^{n-1}\] so in your case, \[6^n~'=6x^{6-1}=6x^5\] so you would be getting 0-6x^5 which is same as just saying -6x^5.

OpenStudy (solomonzelman):

tnx for the medal bro !